OpenStudy (anonymous):

find the distance between the pair of points (3,1)(-2,-2)

OpenStudy (anonymous):



OpenStudy (anonymous):

use the distance formula

jimthompson5910 (jim_thompson5910):

d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((-2-3)^2+(-2-1)^2) d = sqrt((-5)^2+(-3)^2) d = sqrt(25+9) d = sqrt(34) So exact distance between the two points is sqrt(34) units.